OpenStudy (anonymous):

prove by induction. ((n+1)!/(n+1)^(n+1))

OpenStudy (kinggeorge):

Have you done the base case(s) yet?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then (n+2)!/(n+2)^(n+2) , (n+1)!/(n+1)^(n+1) goes to ((n+1)n!)/((n+2)(n+2)^n) < n!/(n+1)^n so (n+1)!/(n+2)^(n+1) < n!/(n+1)^n < n!/n^n so (n+1)!/(n+2)^(n+1) < n!/n^n this is where im stuck

OpenStudy (anonymous):

the first line shold be a < not a ,

OpenStudy (kinggeorge):

Let's see then. We want to show the following.\[\frac{(n+2)!}{(n+2)^{n+2}}<\frac{(n+1)!}{(n+1)^{n+1}}\]We know that \[\frac{(n+1)!}{(n+1)^{n+1}}<\frac{n!}{n^n}\]And we also have\[\frac{(n+2)!}{(n+2)^{n+2}}=\frac{(n+2)(n+1)!}{(n+2)(n+2)^{n+1}}=\frac{(n+1)!}{(n+2)^{n+1}}\]Now, note that \[\frac{1}{(n+2)^{n+1}}<\frac{1}{(n+1)^{n+1}}\]Hence, \[\frac{(n+1)!}{(n+2)^{n+1}}<\frac{(n+1)!}{(n+1)^{n+1}}\]Which is what we wanted to show.

OpenStudy (anonymous):

where did the inductive hypothesis come into play?

OpenStudy (kinggeorge):

Now that I've looked over that, I don't think I ever used it. I don't know why you would even want to use it here. It seems like it would just complicate things.

OpenStudy (anonymous):

yeah, seems so. We dont have to use it?

OpenStudy (kinggeorge):

I've got to say, I don't even know where one could even use the inductive hypothesis. Are you sure the problem is written exactly that way?

OpenStudy (anonymous):

well, this is to show that n!/n^n is always increasing, the book gives a hint that says it suffices to show that ((n+1)!/(n+1)^(n+1))<n!/n^n for all n in N. So I guess I dont have to use induction.

OpenStudy (anonymous):

err decreasing*

OpenStudy (kinggeorge):

In this case, I don't think you need to use induction. However, if it does say to show something for all \(n\in\mathbb{N}\), induction is usually a good bet.

OpenStudy (anonymous):

yeah, thats why I thought it was saying use it. Great ty sir.

OpenStudy (kinggeorge):

You're welcome.